OpenStudy (anonymous):

14. How large must the coefficient of static friction be between the tires and the road if a car is to round a level curve of radius 85 m at a speed of 90 km/h?

OpenStudy (anonymous):

coefficient of static friction=tan(theta)=v^2/r*g. put v=90*5/18=25m/sec. r=85m,and g=9.8m/sec^2 and find coefficient of static friction.

OpenStudy (anonymous):

here the friction supplies the nesessary centripetal force